How would you solve this equation?
The width of a rectangle is 6 cm less than the length. The perimeter is 72 cm. Write and solve an equation to find the width and the length of the rectangle.

Answer by Paul(988)

Let x be the length and let x-6 be the width
2(x+x-6)=72
4x-12=72
4x=84
x=21
21-6=15
Therefore the width is 15 and the length is 21